OpenStudy (anonymous):

Algebra Mixture: How many liters of 7% acid must be mixed with how many liters of 12% acid solution by a chemist to obtain 6 liters of 10% acid?

OpenStudy (anonymous):

Ok so i think this is how you write it but im putting this up to make sure. So first you write .07x+12(6-x)=(6)(0.1) then, when you solve it you get x=2.4

OpenStudy (anonymous):

Anybody know if its right and want to tell me?

OpenStudy (roadjester):

ok, this isn't a mixture...

OpenStudy (anonymous):

It si , it says mixtures on the top of the page... lol

OpenStudy (anonymous):

but whatever you may call it, if you know the answer please help i have like 40 more topics to study and my exam is in 2 days

OpenStudy (anonymous):

Here is how I would think about it. 6 liters of 10% acid is .6 liters of acid in 6 liters total. You have two other solutions your mixing, and you want \(x+y=6\), \(.07x+.12y=.6\). You're solving for \(x\), so put the first equation in terms of \(y\), \(y=6-x\), then substitute: \(.07x+.12(6-x)=.6\). So yeah, that was a long way of saying that your equation was correct.

OpenStudy (anonymous):

oh lol tysm :) you really helped xD

OpenStudy (mertsj):

Let x = amount of 7% solution to add Then 6-x = amount of 12 % solution to add .07x+.12(6-x)=.10(6)

OpenStudy (roadjester):

Ok...the long, roundabout way. lol

OpenStudy (mertsj):

And yes. You did it right.

OpenStudy (anonymous):

Thanks :)

OpenStudy (mertsj):

yw

OpenStudy (roadjester):

I like mertsj answer better, sorry nbouscal

OpenStudy (anonymous):

Both are good and help me alot, thanks to all three of you, ill remember you guys when i take my exam xD lol

OpenStudy (roadjester):

you can just become fans. ☺

OpenStudy (anonymous):

lol i thought i already was for all 3

OpenStudy (mertsj):

I always used to teach my students to write the first equation based on the amount of solution: x+(6-x)=6 A very trivial but useful equation. Not based on that, put in the percents and write the equation based on the amount of solute: .07x+.12(6-x)=.10(6)

OpenStudy (mertsj):

Now based on that, I meant to say.

OpenStudy (roadjester):

Looks like we have a math teacher lol

OpenStudy (anonymous):

xD